Use inductive reasoning to predict the next complete line in the pattern.
44 - 5 = 39
454 - 56 = 398
4554 - 567 = 3987

Answer by Theo(13342)

Use inductive reasoning to predict the next complete line in the pattern.
44 - 5 = 39
454 - 56 = 398
4554 - 567 = 3987

It looks like the the number that is being subtracted is being multiplied by 10 and then having the next higher digit added to that.

5 * 10 = 50 + 6 = 56

56 * 10 = 560 + 7 = 567

If that's true, then the next number to be subtracted will be:

567 * 10 = 5670 + 8 = 5678 *****

It looks like the the number that is being added is being multiplied by 10 and then having 14 added to it.

44 * 10 = 440 + 14 = 454

454 * 10 = 4540 + 14 = 4554

If that's true, then the next number to be added will be:

4554 * 10 = 45540 + 14 = 45554 *****

If all of that is correct, then the next complete line will be:

45554 - 5678 = 39876

The next complete line after that will be:

455554 - 56789 = 398765

etc...